OpenStudy (anonymous):

You are standing in line to see a movie. Five more people are ahead of you in line than are behind you. Four times as many people are in line as the number of people who are behind you. A: How many people are ahead of you in line? B: How many people are in line including you?

OpenStudy (anonymous):

Is it 9 or 5? please help me!!!!!!!!!

OpenStudy (jojokiw3):

Kk so let's say that x = people in front and y = people behind you. x = y + 5 Since x = y + 5 and the total number of people = 4y, the equation is: y+y+5+1 = 4y Now you add everything together, resulting in 2y + 6 = 4y. Add -2y to both sides, and you get 6 = 4y - 2y. 6=2y. 6/2 = y y = 3. x = 3 + 5 x = 8. y = people behind you. x = people in front of you. you = 1. 1 + 8 + 3 = 12. 4y=12 A. 8 B. 12

OpenStudy (anonymous):

OMG thank you that was really helpful!!!!!!